The Phi Nancy Company will lend you $20,000 for your new car purchase. Phi Nancy requires that you pay that load back in monthly payments of $950 cash, for a total of thirty-six monthly payments.

What is the annual percentage rate (APR) and the effective annual rate (EAT) that Phi Nancy require on this loan?
Answer #1

Given,

Loan amount (PV) = $ 20,000

Number of periods(nper) = 36

Monthly amount(PMT) = 950

Rate(rate) = =RATE(36,-950,20000 RATE(nper, pmt, pv, [fv], [type]. = 3.25% per month

Annual Percentage Rate = 3.25% * 12 = 38.95% pa

EAR = (1+ APR/12)^12 -1 ( assuming that the interest rate is compounded monthly)

EAR = (1+ 38.95%/12)^12 - 1 = 46.72%